### v2

# Temporal Misalignment Attacks against Multimodal Perception in Autonomous Driving

Abstract
Multimodal fusion (MMF) plays a critical role in the perception of autonomous driving, which primarily fuses camera and LiDAR streams for a comprehensive and efficient scene understanding. However, its strict reliance on precise temporal synchronization exposes it to new vulnerabilities. In this paper, we introduce DejaVu, an attack that exploits the in-vehicular network and induces delays across sensor streams to create subtle temporal misalignments, severely degrading downstream MMF-based perception tasks. Our comprehensive attack analysis across different models and datasets reveals the sensors’ task-specific imbalanced sensitivities: object detection is overly dependent on LiDAR inputs, while object tracking is highly reliant on the camera inputs. Consequently, with a single-frame LiDAR delay, an attacker can reduce the car detection mAP by up to 88.5%, while with a three-frame camera delay, multiple object tracking accuracy (MOTA) for car drops by 73%. We further demonstrated two attack scenarios using an automotive Ethernet testbed for hardware-in-the-loop validation and the Autoware stack for end-to-end AD simulation, demonstrating the feasibility of the DejaVu attack and its severe impact, such as collisions and phantom braking.
I Introduction
Autonomous driving (AD) is designed to navigate and interact with complex environments—a capability fundamentally reliant on a comprehensive understanding of its surroundings. The adoption of heterogeneous sensors, such as camera, LiDAR, and radar, that capture data from different modalities allows an accurate perception with enhanced accuracy and robustness [1]. Each individual modality has unique strengths; for example, cameras capture rich semantic details, LiDAR provides accurate depth measurements, and radar particularly excels in detecting speed, even in adverse weather conditions [2]. However, these sensors also face inherent limitations, such as a camera’s sensitivity to lighting variations, LiDAR’s lack of texture information, and radar’s sparsity, which can compromise performance when used independently [3, 4, 5]. Multimodal fusion (MMF)—the process of integrating multiple unimodal sensor data into a single and comprehensive representation—compensates for such individual sensor weaknesses, ensures accurate and robust perception, and efficient downstream tasks in AD [6, 7].
MMF remains a fundamental research challenge, primarily due to the heterogeneity across sensing modalities, including differences in data formats, spatial resolutions, and temporal characteristics [8, 9]. However, the robustness of temporal alignment—ensuring that all sensor data being fused corresponds to the same point in time—has received comparatively limited attention [10], despite being the fundamental enabler of MMF. Particularly, in AD, temporal alignment is a fundamental prerequisite to enable real-time perception, which directly informs decision-making and control. Misalignments, caused, for instance, by delays in specific sensor streams, can degrade perception performance, leading to object misdetection, localization errors, and scene misinterpretation. These inaccuracies can propagate downstream, affecting core functions such as control, maneuver planning, and safety interventions, ultimately compromising the reliability and safety of the vehicle [11]. Moreover, temporal alignment in MMF for AD remains a complex problem with the following challenges.
Figure 1: Impact of DejaVu attack on 3D object detection. (a-b) show two benign scenarios without any temporal misalignment, and hence, accurate object detection. (c-d) Illustrate different temporal alignments between camera and LiDAR inputs and highlight how delayed sensor data can lead to incorrect detections, either by detecting non-existent objects (false positives) or missing present ones (false negatives). In (c), the MMF prioritizes the (delayed) LiDAR data and predicts three objects, including a pedestrian who is not present in the current camera view, resulting in seemingly accurate results from LiDAR’s perspective but a false detection from the camera’s perspective. However, in (d), the MMF still prioritizes the (updated) LiDAR data and predicts two objects, excluding the pedestrian who is present in the current camera view, resulting in a missed detection from the camera’s perspective. In both temporal misalignment attack cases, MMF biased its fusion toward LiDAR, failing to account for semantic discrepancies in the camera modality.
Clock synchronization.  To ensure temporally aligned perception in AD, sensor data from all the modalities must be timestamped relative to a common global clock with minimal drift. In practice, the electronic controller units (ECUs) hosting the sensors must be tightly synchronized—often at sub-microsecond precision required by AUTOSAR [12])—to enable accurate fusion and downstream reasoning. To achieve that, automotive Ethernet (AE) equipped with Time-Sensitive Networking (TSN) has become the de facto backbone of modern in-vehicle networks. TSN leverages the generalized Precision Time Protocol (gPTP, IEEE 802.1AS) to maintain global clock synchronization among distributed ECUs [13, 14]. Consequently, a core security assumption is: $\scriptstyle A_{1}$ The gPTP-based synchronization infrastructure is secure and maintains consistent global time across all ECUs. However, despite its widespread adoption in AE, gPTP was primarily designed to provide deterministic time synchronization—not to operate under adversarial conditions. Hence, gPTP lacks built-in security mechanisms, particularly cryptographic authentication, leaving it vulnerable to attacks such as grandmaster spoofing, delay injection, replay, and false time advertisement [15, 16, 17]. Such attacks can compromise the temporal alignment of sensor fusion, undermining the integrity of MMF-based perception, without any direct alteration to the sensor data or sensor ECU itself.
Timestamp Integrity.  Even when sensors are nominally synchronized, they often capture data at slightly different times due to variations in sampling rates caused by inherent hardware limitations [10]. Consequently, the timestamps attached to sensor messages become critical in multimodal fusion pipelines, as they serve as the primary reference for aligning asynchronous data streams by identifying the most temporally proximal pairs from buffered queues. This reliance gives rise to a fundamental security assumption: $\scriptstyle A_{2}$ Sensor ECUs are trusted entities that always provide accurate timestamps. However, this assumption can be violated. In vehicular systems, sensor ECUs may be compromised through remote exploits of unpatched software vulnerabilities [18, 19, 20], insecure Over-The-Air (OTA) updates [21, 22], or direct physical access via the OBD-II port [23, 24, 18]. Once compromised, an attacker-controlled ECU can inject messages containing legitimate sensor data but with forged timestamps. Since the fusion ECU uses timestamps as the basis for temporal alignment, this can result in the selection of data pairs that appear temporally aligned but are actually misaligned, thereby degrading the integrity of the sensor fusion process.
Middleware Integrity.  Sensor data exchange in production-grade AD (such as Autoware [^1] https://github.com/autowarefoundation/autoware) is often facilitated by robotic middleware frameworks, such as robot operating system (ROS [^2] https://github.com/ros2/ros2), where the data distribution service (DDS) serves as the underlying communication backbone. DDS enables a real-time and scalable publish-subscribe communication model for data sharing among distributed ECUs, making it a widely adopted solution for managing the high-bandwidth, low-latency communication demands of multimodal perception and control pipelines. Consequently, a third foundational assumption arises: $\scriptstyle A_{3}$ The ROS-based DDS infrastructure is secure and ensures the integrity, authenticity, and freshness of shared data. However, the design of ROS prioritizes performance and scalability over robust security guarantees. Since ROS does not enforce strong authentication, encryption, or source verification by default, ROS is vulnerable to a wide range of network-level and application-level attacks, including message spoofing, replay, and impersonation attacks [25]. As a result, an attacker with access to the ROS communication graph, for example, can impersonate legitimate nodes (e.g., a LiDAR publisher) and can publish fabricated sensor messages with targeted timestamps (i.e., replay attacks), which can be considered by the fusion ECU and mislead downstream perception tasks.
Moreover, prior work has demonstrated that existing MMF-based perception systems in AD lack temporal robustness—the ability to maintain reliable outputs in the presence of temporal inconsistencies—making them vulnerable even to benign, non-malicious misalignments [10, 26, 27]. In particular, perception pipelines have been shown to degrade significantly under small delays in just one modality. Existing studies, however, are limited in both scope and threat modeling: they primarily focus on random, system-induced delays and evaluate only a narrow subset of perception tasks—most commonly 3D object detection, without realizing them on any end-to-end AD software stacks. To investigate the true fragility of MMF-based perception under adversarial conditions, we present a comprehensive study of temporal misalignment attacks, which we term DejaVu. These attacks are realized by violating one or more of the core trust assumptions. Unlike prior work, we target both 3D object detection and multi-object tracking (MOT), and evaluate the effects of different delay distributions across multiple perception models and datasets. Fig. 1 illustrates how temporal delays due to DejaVu attacks in one of the modalities can degrade the performance of the MVXNet model [28]—a MMF-based 3D object detection model, potentially leading to unsafe driving conditions in AD.
In summary, we make the following key contributions:
  • We propose DejaVu, a temporal misalignment attack against MMF in AD that exploits vulnerabilities of in-vehicle networks and the fragility of multimodal fusion by selectively delaying sensor streams to disrupt perception in safety-critical tasks.
  • We conduct a comprehensive empirical evaluation of DejaVu on state-of-the-art 3D object detection models (MVXNet [28], BEVFusion [29]) using the KITTI [30] and nuScenes [31] datasets, and a multi-object tracking model (MMF-JDT [32]) under various misalignment scenarios using the KITTI [30] dataset. Our findings reveal distinct modality-specific vulnerabilities: object detectors are highly sensitive to LiDAR delays, while the tracking model is significantly impacted by camera timing disruptions. A single-frame LiDAR delay reduces 3D detection mAP by up to 88.5%, and a three-frame camera delay drops multiple object tracking accuracy (MOTA) by 73%.
  • To further validate our findings in a realistic autonomous driving setting, we built an automotive Ethernet testbed that models the sensor data acquisition and fusion pipeline. Using this platform, we implemented the DejaVu attack by violating $\scriptstyle A_{2}$, demonstrating its feasibility in a hardware-in-the-loop environment. Additionally, to assess the end-to-end consequences beyond perception—specifically on planning and control—we integrated DejaVu into Autoware, a production-grade, full-stack autonomous driving simulator, by breaking $\scriptstyle A_{3}$. In both environments, our experiments show that DejaVu is highly practical and can result in severe safety violations, including direct collisions and phantom braking events.
The remainder of the paper is organized as follows. Section II reviews related work; Section III outlines the threat model and proposes DejaVu attack; Section IV describes datasets and evaluation settings; Section V reports the evaluation results; Section VI reports the discussion and potential defenses; and Section VII concludes the paper.
II Related Work
The fundamental research direction has been in the direction of spoofing sensors from a single modality, such as LiDAR [33, 34, 35] and camera [36] through different means of physical perturbation. Moreover, advanced attacks have demonstrated to even compromise multiple modalities together [37, 38, 39]. These sensor spoofing attacks demonstrate that simply having multiple sensors is not sufficient. With carefully constructed inputs, an adversary can simultaneously mislead camera and LiDAR sensors, defeating the very redundancy meant to ensure safety.
Unlike sensor spoofing attacks, time delay attacks have not been extensively studied within AD. They have created significant attention in other cyber-physical systems (CPS) domains, such as power systems [40, 41], wireless networks [42], unmanned aerial vehicles (UAVs) [43, 44, 15], and time-sensitive networks [45]. Software timing interference is also exploited to cause system destabilization in CPS [46]. Moreover, multimodal temporal misalignment in sensor fusion has been shown to degrade the accuracy of simultaneous localization and mapping (SLAM) [47], which was limited only to the fusion between IMU and camera data. Contrary to the existing works, we comprehensively study the impact of a temporal misalignment attack on task-agnostic MMF-based perception.
Figure 2: Overview of the proposed system modeling with DejaVu attack.
In the realm of AD security, various defense mechanisms have been proposed to counteract sensor spoofing [48] and multimodal fusion attacks. These defenses can be broadly categorized into spatiotemporal consistency checks, specification-aware recovery strategies, and hardware-based techniques [49]. PercepGuard [50] detects misclassification attacks by enforcing consistency between object tracks and class labels, but it does not examine the temporal validity of sensor readings and thus cannot detect replayed LiDAR scans whose semantic trajectories remain plausible. Connecting the Dots [51] employs class‐specific autoencoders to uncover context violations introduced by adversarial perturbations, yet time‐shifted data aligns perfectly with learned scene co‐occurrence statistics and evades its checks. PhyScout [52] formalizes cross‐modal conflict detection to identify gross spoofing, but subtle timestamp manipulations within the synchronizer’s tolerance window introduce no overt spatial or modality discrepancies. These approaches overlook timestamp integrity and data freshness as a security property.
III Temporal Misalignment Attack : DejaVu
This section elaborates on the system model, the threat model, and the details of the proposed temporal misalignment attacks. Table I contains a summary of notations used throughout the paper.
TABLE I: Summary of Notations
Symbol Description
$S_{i}$ $i$-th unimodal sensor
$m$ Total number of sensor modalities
$(x_{i}^{(t_{\text{act}})},t_{\text{pre}})$ Data packet from $S_{i}$ with capture and transmission times
$x_{i}^{(t_{\text{act}})}$ Sensor reading from $S_{i}$ captured at time $t_{\text{act}}$
$t_{\text{pre}}$ Time when $x_{i}^{(t_{\text{act}})}$ is made available or transmitted
$E_{i}$ Encoder for modality $i$ (unimodal feature extractor)
$f_{i}^{(t_{i})}$ Feature representation of input $x_{i}^{(t_{i})}$
$\mathcal{F}_{\theta}$ Multimodal fusion network
$f_{mm}^{(t_{\text{sys}})}$ Multimodal fused representation at system time $t_{\text{sys}}$
$\mathcal{H}_{\theta}$ Perception head (task-specific prediction layer)
$\delta_{i}^{\text{syn}}$ Clock drift (nominal temporal misalignment) for sensor $i$
$\delta_{i}^{\text{com}}$ Communication latency (nominal) for sensor $i$
$\delta_{i}^{\text{mal}}$ Added malicious delay for sensor $i$
$k$ Delay control threshold or parameter
${\mathcal{Y}}^{(t_{\text{sys}})}$ Predicted output (benign) at time $t_{\text{sys}}$
$\tilde{\mathcal{Y}}^{(t_{\text{sys}})}$ Predicted output (under attack or corrupted) at $t_{\text{sys}}$
III-A System Modeling
As shown in Fig. 2, we consider an MMF-based perception system with $m$ sensors denoted as $S_{1},S_{2},\ldots,S_{m}$. At any time $t_{\text{act}}$, the transmitted sensor data from $S_{i}$ is represented by $(x_{i}^{(t_{\text{act}})},t_{\text{pre}})$.
Definition 1 ($x_{i}^{(t_{\text{act}})},t_{\text{pre}}$).
The tuple $(x_{i}^{(t_{\text{act}})},t_{\text{pre}})$ represents a sensor data packet where $x_{i}^{(t_{\text{act}})}$ denotes the data “actually” captured at time $t_{\text{act}}$ (as indicated by the superscript), and $t_{\text{pre}}$ is the associated timestamp metadata, indicating the time at which the sample is “presumed” to have been captured. Discrepancies between $t_{\text{act}}$ and $t_{\text{pre}}$ suggest temporal misalignment.
In MMF-based techniques timestamp $t_{\text{pre}}$ is used to ensure proper synchronization before fusing them and under a benign scenario at time $t_{i}$, the sensor data packet data can be expressed as $(x_{i}^{(t_{i})},t_{i})$, where $t_{i}=t_{\text{act}}=t_{\text{pre}}$. As each of the heterogeneous sensors generates data samples with a unique format and dimension, to ensure an effective sensor fusion, a modality-specific unimodal feature encoder (UFE) $E_{i}$ converts the raw data $x_{i}^{(t_{i})}$ to an intermediate representation $f_{i}^{(t_{i})}=E_{i}(x_{i}^{(t_{i})})$. Such encoders map the raw data to a similar feature format, which allows different sensor fusion techniques, such as straightforward operations like concatenation, merging, average polling, etc., or even more advanced techniques such as tensor-based fusion [53]. Assume that sensor fusion is performed at a global time $t_{\text{sys}}$, using a fusion function denoted by $\mathcal{F}_{\theta}$. This process yields a fused representation $f_{mm}^{(t_{\text{sys}})}$, computed as follows:
$f_{mm}^{(t_{\text{sys}})}=\mathcal{F}_{\theta}\big(f_{1}^{(t_{1})},f_{2}^{(t_{2})},\ldots,f_{m}^{(t_{m})}\big)$
Finally, the corresponding perception head for the downstream tasks $\mathcal{H}_{\theta}$ predicts the perception results from the fused representation as:
$\mathcal{Y}^{(t_{\text{sys}})})=\mathcal{H}_{\theta}\big(f_{mm}^{(t_{\text{sys}})}\big)$
As $t_{i}$ is the actual time when the sensor $S_{i}$’s data was captured, and $t_{\text{sys}}$ is the time when the sensor data are synchronized and fused at the central ECU, under normal conditions, the $t_{i}$ can be expressed as:
$t_{\text{sys}}=t_{i}+\delta_{i}^{\text{syn}}+\delta_{i}^{\text{com}},\quad\forall i\in\{1,\dots,m\}$ (1)
where $\delta_{i}^{\text{syn}}$ represents the nominal temporal misalignment (clock drift) from the global time due to the clock drift of $S_{i}$. This delay is sensor-specific and under a perfect clock synchronization, it becomes zero ($\delta_{i}^{\text{syn}}\approx 0$). On the other hand, $\delta_{i}^{\text{com}}$ is the communication latency, which indicates the time needed for the data to reach the sensor fusion module. With fast communication infrastructure (AE with TSN) with efficient middleware, this can be eliminated ($\delta_{i}^{\text{com}}\approx 0$). Therefore, under an ideal condition (such as $\scriptstyle A_{1}$$\scriptstyle A_{3}$), $t_{\text{sys}}\approx t_{1}\approx t_{2}\approx\dots\approx t_{m}$ for $m$ sensors, ensuring that data from all sensors corresponds to the same system time $t_{\text{sys}}$.
III-B Threat Model
III-B1 Attacker Capabilities
In the threat model with an adversary capable of gaining access to the in-vehicle network through one or more of the following realistic entry points [54]: Physical Access: The attacker connects to the vehicle’s OBD-II port or directly to Ethernet/CAN interfaces, either through malicious maintenance personnel or via physical compromise [18]. Remote Exploitation: The attacker exploits vulnerabilities in externally exposed interfaces, such as telematics units, infotainment systems, or over-the-air (OTA) update mechanisms [18, 20, 22]. Supply Chain Attacks: The attacker implants malicious code or hardware during manufacturing, allowing persistent access post-deployment [55]. Once access is established, the attacker can monitor, intercept, and inject messages on the in-vehicle AE backbone and associated sub-networks. This enables manipulation of time-sensitive communications, particularly those involved in clock synchronization and forged data/timestamp propagation. We assume the attacker operates under any of these three distinct capabilities.
$\scriptstyle C_{1}$
Disruption of Clock Synchronization.  With this capability, the adversary targets the clock synchronization mechanism in AE—specifically, the gPTP. Rather than altering timestamps directly, the attacker compromises the synchronization process itself, thereby inducing actual temporal misalignment between sensor streams and other sub-networks. This can be accomplished by impersonating the grandmaster clock or tampering with synchronization messages via selective delay, replay, or man-in-the-middle attacks [56, 57, 58]. Although individual timestamps remain unaltered, they no longer correspond to a consistent global time due to induced drift, impairing the performance of the downstream perception task.
$\scriptstyle C_{2}$
Manipulation of Timestamp Integrity.  In this scenario, the attacker is capable of preserving the actual timing of data transmission but alters the timestamps embedded in transmitted packets [54]. This creates a seemingly temporally misaligned messages in the data queue from the timestamps perspective while the data contents are, in fact, temporally aligned. Consequently, while the data synchronizer [^3] TimeSynchronizer and ApproximateTimeSynchronizer are commonly used message filtering utilities in ROS2 that align multiple sensor message streams based on their timestamps. While TimeSynchronizer performs strict timestamp matching, ApproximateTimeSynchronizer allows messages with slight temporal differences—within a specified tolerance window—to be synchronized. finds the seemingly aligned inputs, in reality, it finds actual temporal misaligned pairs.
$\scriptstyle C_{3}$
Impersonation of a Legitimate Node in ROS2.  In this scenario, the attacker is assumed to be a participant in the ROS2 network. As the default ROS2 implementation lacks built-in security mechanisms—a configuration commonly adopted in industry-grade AD software stacks, including Autoware—the attacker can instantiate a malicious node that impersonates a legitimate sensor publisher. By subscribing to target sensing topics [^4] Representative topic names from Autoware are /sensing/camera/traffic_light/image_raw, /sensing/lidar/top/pointcloud_raw, etc., the attacker gains access to the data stream and records historical sensor messages. The malicious node then re-publishes previously captured messages with updated, genuine-looking timestamps, potentially while the original publisher is still active. By repeatedly injecting such delayed-but-valid messages, the attacker can pollute the input queue of time-based synchronizers, increasing the likelihood that a forged message is selected during the fusion process.
Thus, we consider an attacker equipped with either of these capabilities or even multiple of them to launch the DejaVu attack, thereby causing temporal misalignment and compromising perception integrity.
III-B2 Attackers Objective
The adversary’s primary objective is to compromise the integrity and reliability of the sensor fusion by introducing deliberate temporal misalignment in one or more sensors. The attacker aims to disrupt the coherent integration of multi-modal sensor data, thereby inducing errors in perception. The resultant temporal inconsistencies can lead to incorrect perception results, particularly in safety-critical applications such as AD. Hence, the attacker’s key objective is to disrupt the perception to deteriorate the downstream tasks and, eventually, to compromise the safety and integrity of the control decision.
III-C Proposed DejaVu Attack
In this part, we elaborate on our proposed temporal misalignment attack, named DejaVu attack, where an adversary maliciously introduces a timing delay $\delta_{i}^{\text{mal}}$ to one or more sensors with a goal of creating misaligned sensor fusion and creating a false perception of the surroundings. Under DejaVu attack, the attacker compromises the network to add an additional malicious delay of $\delta_{i}^{\text{mal}}$. Hence, the attacker transmits a compromised data packet $\big(\tilde{x}_{i}^{(\tilde{t}_{i})},t_{i}\big)$ to the fusion ECU with the outdated semantic content $\tilde{x}_{i}^{(\tilde{t}_{i})}$ which was captured at global time $\tilde{t}_{i}$ instead of $t_{i}$, but with the updated time stamp $t_{i}$. In this case, $t_{i}=\tilde{t}_{i}+\delta_{i}^{\text{mal}}$, and the attacker delayed that transmission by $\delta_{i}^{\text{mal}}$ either through (such as $\scriptstyle C_{1}$$\scriptstyle C_{3}$). Therefore, $t_{\text{sys}}$ can be expressed as:
$t_{\text{sys}}=\underbrace{\tilde{t}_{i}+\delta_{i}^{\text{mal}}}_{\displaystyle{t_{i}}}+\delta_{i}^{\text{syn}}+\delta_{i}^{\text{com}}$ (2)
III-C1 Distribution of Delay $\delta_{i}^{\text{mal}}$
The malicious delay $\delta_{i}^{\text{mal}}$ can vary over time–depending on the attacker’s intent, and be crafted to cause maximum disruption in sensor fusion. In a benign scenario, there will be no temporal misalignment, and sensor data will be received and processed in a timely manner. However, the attacker can launch different forms of DejaVu attack by controlling $\delta_{i}^{\text{mal}}$ in different ways. For instance, in Table II, we provide two possible DejaVu attacks scenarios against $S_{i}$, which are elaborated below:
TABLE II: Two Types of DejaVu Attack Strategy Against $S_{i}$
Attack Name Attack Type Delay Distribution $\delta_{i}^{\text{mal}}$
Constant Delay Constant Constant, $k$
Random Delay Random $\text{Uniform}(0,k)$
Constant Delay
This attack strategy introduces a fixed delay of $k$ frames in the sensor data stream. The attacker can achieve this by creating desynchronization among the clocks, tampering with sensor timestamps, or delaying data at the communication layer. Impact: The fusion model still receives temporally consistent data but with a lag in the target modality. This can lead to delayed (misaligned) perception (i.e., missing or shifted bounding boxes) in critical applications like object detection, where real-time perception is essential.
Random Delay
In this attack strategy, each frame experiences a different delay, randomly sampled from the range $[0,k]$. This strategy not only disrupts the real-time requirement but also disrupts the temporal sequence of sensor data, which is crucial for object tracking. Impact: The fusion system struggles to maintain the proper ordering of sensor inputs, leading to degraded perception accuracy. This will cause erratic behavior in time-sensitive sequential applications, particularly for object tracking and autonomous navigation.
III-C2 Impact of the DejaVu Attack
Based on the attacker’s capabilities, she can compromise either only one modality, which we define as an unimodal DejaVu attack (Uni-DejaVu), or she can compromise multiple modalities together, defined as a multimodal DejaVu attack (Mul-DejaVu). Under the Uni-DejaVu attack with compromised, where only sensor $S_{1}$ is compromised, the compromised perception result $\tilde{\mathcal{Y}}^{(t_{\text{sys}})}$ at time $t_{\text{sys}}$ can be expressed as:
$\tilde{\mathcal{Y}}^{(t_{\text{sys}})}=\mathcal{H}_{\theta}\big(\mathcal{F}_{\theta}\big(\tilde{f}_{1}^{(\tilde{t}_{1})},f_{2}^{({t}_{2})},\dots,f_{m}^{({t}_{m})}\big)\big)$ (3)
Similarly, under the Mul-DejaVu attack with all the sensors compromised, the compromised perception result $\tilde{\mathcal{P}}(t_{\text{sys}})$ at time $t_{\text{sys}}$ can be expressed as:
$\tilde{\mathcal{Y}}^{(t_{\text{sys}})}=\mathcal{H}_{\theta}\big(\mathcal{F}_{\theta}\big(\tilde{f}_{1}^{(\tilde{t}_{1})},\tilde{f}_{2}^{(\tilde{t}_{2})},\dots,\tilde{f}_{m}^{(\tilde{t}_{m})}\big)\big)$ (4)
Under the DejaVu attack, the system will perceive the surroundings in a different way than that of the benign case $\tilde{\mathcal{Y}}^{(t_{\text{sys}})}\neq{\mathcal{Y}}^{(t_{\text{sys}})}$, which can lead to erroneous vehicle control decisions.
IV Experimental Settings
This section describes the experimental setup to assess the DejaVu attack across different MMF models and datasets.
IV-A Datasets
We evaluate our approach on two widely used multimodal AD datasets for 3D object detection and multi-object tracking: KITTI Tracking and NuScenes.
[…]
NuScenes Dataset.  The NuScenes Dataset [31] was collected in Boston (USA) and Singapore, focusing on dense urban traffic under diverse conditions. It provides six surround-view RGB cameras, a Velodyne HDL-32E LiDAR (20 Hz), and five radars, with all annotations sampled at 2 Hz. The dataset consists of 1000 driving scenes, each 20 seconds long, with 3D bounding boxes and tracking IDs for 23 classes, including vehicles, pedestrians, bicycles, and traffic barriers.
Figure 3: Uni-DejaVu attack impacts on 3D object detection performance of MVXNet on KITTI dataset for different object classes.
Figure 4: Mul-DejaVu attack impacts on 3D object detection performance of MVXNet on KITTI dataset for different object classes.
Figure 5: Uni-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for different object classes.
IV-B Models
To systematically evaluate the effect of DejaVu on MMF with different downstream tasks, we consider the following models:
3D Object Detection: We evaluate DejaVu on two representative MMF-based 3D object detection models. i) MVXNet [28], trained on the KITTI dataset, is an early fusion-based architecture that projects LiDAR point clouds into pseudo-image space and fuses them with camera image features at the voxel level. ii) BEVFusion [29] is a more advanced architecture, trained on the NuScenes dataset, that unifies multi-modal sensor inputs in the bird’s eye view (BEV) representation space. BEVFusion is widely used in industry-grade AD software stacks, including Autoware.
Multi-Object Tracking (MOT): For evaluating tracking performance under DejaVu, we consider MMF-JDT [32], trained on the KITTI tracking dataset, is a joint detection and tracking model that incorporates early and mid-level fusion strategies to align image and point cloud features for improved object association over time.
Figure 6: Mul-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for different object classes.
IV-C Attack Settings
To evaluate the impact of DejaVu attack on MMF models, we introduce controlled delays in one or both sensor modalities (camera and LiDAR), as mentioned in Table II, and analyze the corresponding degradation in model performance. We systematically assess how different degrees and types of temporal misalignment affect the 3D object detection and multi-object tracking. We introduce both constant and random delays in one or both modalities, as defined in Table II. Specifically, we use a delay parameter $k\in\{0,1,2,3,4,5\}$, where $k=0$ represents a perfectly synchronized sensor, and $k=5$ denotes a maximum delay of five frames for the affected modality. We consider both Uni-DejaVu and Mul-DejaVu attacks. In Uni-DejaVu attack, delay is introduced in either the camera or the LiDAR input while keeping the other modality synchronized. In Mul-DejaVu attacks, both camera and LiDAR streams are delayed independently, leading to varying degrees of temporal misalignment. We use the pretrained weights for the target model provided with the official implementations. We focus exclusively on attacking the test dataset by applying the defined temporal delays.
IV-D Evaluation Metrics
To assess the impact of DejaVu, we analyze the model performance using task-specific evaluation metrics. For 3D object detection, we evaluate the models using mean average precision (mAP), which quantifies the accuracy of detected objects, as well as nuScenes Detection Score (NDS), particularly for the NuScenes dataset. For MOT, we use standard tracking metrics such as higher order tracking accuracy (HOTA), detection accuracy (DetA), association accuracy (AssA), multiple object tracking accuracy (MOTA), and Identity Switches (IDSW), which measure the effectiveness of object association across frames.
IV-E Software Implementation
We implement and evaluate DejaVu using Python 3.8 and PyTorch, utilizing open-source frameworks including MMDetection3D [59] and OpenPCDet [60]. Experiments were conducted on a server running Ubuntu 20.04.6 LTS with an Intel Xeon Gold 5520 (16 cores, 2.20GHz), 128GB RAM, and three NVIDIA RTX 6000 Ada GPUs.
V Evaluation Results
This section presents the DejaVu attack impact on different MMF models and datasets, and discusses the key findings from the evaluation.
V-A Impact of DejaVu Attack on 3D Object Detection
We investigate the effectiveness of the proposed DejaVu attack on multimodal 3D object detection using MVXNet and BEVFusion, evaluated on the KITTI and nuScenes datasets, respectively. We analyze the detection performance of 3D object detection across different object classes under varying levels of unimodal and multimodal sensor delay. Object detection models do not process sequential information; instead, their performance is affected only by frame-wise delays in each modality at any particular time, whether the delays are constant or random. Therefore, for simplicity and consistency, we only analyze the attack under the constant delay setting.
Figure 7: Overall Mul-DejaVu attack impacts on 3D object detection performance of BEVFusion on nuScenes dataset for all the object classes regarding i) NDS, ii) mAP.
(a) Constant Delay Scenario
(b) Random Delay Scenario
Figure 8: DejaVu attack impacts on multi-object (car) tracking performance of MMF-JDT on KITTI tracking dataset with respect to different metrics.
V-A1 MVXNet on KITTI Dataset
Fig. 3 presents the 3D object detection performance of MVXNet under Uni-DejaVu attacks, where either the camera or LiDAR input is delayed independently. Under benign (zero-delay) conditions, MVXNet achieves strong mAP across most object classes: approximately 84.1 for cars, 87.6 for pedestrians, and 73.2 for cyclists. The left plot shows that delaying the camera input alone–under Uni-DejaVu camera attacks—has minimal effect on performance across all classes, with nearly constant mAPs. In contrast, the right plot highlights the model’s high sensitivity to LiDAR delays: a 1-frame delay causes the car mAP to collapse from 84.1 to 9.7 ($\downarrow$88.5%), pedestrian mAP from 87.6 to 34.2 ($\downarrow$60.9%), and cyclist mAP from 73.2 to 32.9 ($\downarrow$55.1%), with further degradation as delay increases. This indicates that LiDAR data is significantly more critical than camera input in MVXNet’s perception pipeline; hence, MVXNet’s high vulnerability against Uni-DejaVu attack against LiDAR. However, Fig. 4 shows the mAP heatmaps across combinations of camera and LiDAR delays under Mul-DejaVu attacks. Although the 1-frame LiDAR delay drops mAP from 55.1% to 88.5% for different objects, camera delay has almost no effect, indicating the dominance of LiDAR in MVXNet.
{keyfinding}
Key Findings.  MVXNet heavily depends on LiDAR input for 3D object detection. Camera delay, under both Uni- or Mul-DejaVu attack, has almost no effect, but even minor LiDAR misalignment leads to severe performance degradation. When both modalities are delayed, the impact is still only dominated by the LiDAR stream.
V-A2 BEVFusion on nuScenes Dataset
Fig. 5 shows BEVFusion’s 3D detection performance under Uni-DejaVu attacks. With no delay, the model achieves high mAP across most classes—approximately 88 for car and pedestrian, and slightly lower for the rest. With a 1-frame camera delay, mAP drops slightly across all classes (for instance, car mAP drops by 7.4%), but remains stable with further delays, showing BEVFusion is slightly vulnerable against Uni-DejaVu camera attacks. In contrast, introducing a 1-frame LiDAR delay results in a substantial reduction in mAP for specific object classes, with performance dropping by 65.6% for cars (from 88.8 to 30.5), 35.1% for pedestrians (from 88.1 to 57.2), and 38.6% for motorcycles (from 75.9 to 46.6). Although performance remains stable with further LiDAR delays, this finding shows BEVFusion’s higher vulnerability against Uni-DejaVu LiDAR attacks. Similarly, Fig. 6 illustrates BEVFusion’s performance under Mul-DejaVu attacks for all the objects. Although delaying the camera alone has a minimal impact on performance, combining this with a delay in the LiDAR stream causes a significant drop in mAP. For instance, a 1-frame delay in the camera or LiDAR stream reduces car mAP by 7.4% and 65.5%, respectively. However, when both modalities are delayed simultaneously by one frame, the mAP drops dramatically by 89.2% (from 88.8 to 9.6), highlighting a substantial compounding effect. This trend is consistently observed across all object classes. Fig. 7 further shows the impact of Mul-DejaVu on the performance of BEVFusion on the nuScenes dataset regarding average mAP and NDS score.
{keyfinding}
Key Findings.  BEVFusion is slightly affected by camera delays but is highly affected by LiDAR delays, further underscoring LiDAR’s dominating role in 3D object detection. However, the impact becomes significant if both sensors are delayed simultaneously, even just by one frame.
Figure 9: Schematic diagram of hardware-in-the-loop Automotive Ethernet testbed.
Figure 10: ROS System time vs (top) Timestamp of Camera and LiDAR messages, and (bottom) Message content of Camera and LiDAR messages with
Figure 11: Demonstration of replay attacks where malicious LiDAR messages closely follow genuine ones, increasing synchronization likelihood.
(a) False negative scenario.
(b) False positive scenario.
Figure 12: Impact of delayed yet valid LiDAR data on Autoware. a) The ego misses the oncoming truck, causing a head-on collision, and b) the Ego vehicle brakes for a non-existent object from delayed data.
V-B Impact of DejaVu Attack on Multi Object Tracking
We investigate the effectiveness of the proposed DejaVu attack on MOT algorithm using MMF-JDT on the KITTI dataset. We analyze the tracking performance for cars under varying levels of delays under Uni- and Mul-DejaVu attack scenarios.
V-B1 MMF-JDT on KITTI Dataset
This part studies the impact of DejaVu attacks on MMF-JDT evaluated on the KITTI tracking dataset. For this evaluation, we consider both types of delays: constant (Fig. 8a) and random (Fig. 8b). Across both attack scenarios, the tracking performance declines as camera delay increases. Along with other tracking metrics, MOTA and IDSW suffer noticeable degradation as camera delays increase under Uni-DejaVu attacks. For instance, IDSW—a metric that captures identity switches and ideally should be low, increases dramatically under increasing camera delays, underscoring the disruption in tracking consistency caused by DejaVu attacks.
Furthermore, the values of different metrics across the heatmaps of Fig. 8a suggest that while camera delay under Uni-DejaVu deteriorates the performance, delaying both the camera and LiDAR by exactly the same delay (i.e., constant delay scenario) under Mul-DejaVu attack diminishes the attack impact and mostly retains the performance. On the other hand, heatmaps in Fig. 8b show that Mul-DejaVu attacks with random delays remain effective as different delays in both modalities break the sequence, making object tracking considerably more difficult. For instance, under the Mul-DejaVu constant attack scenario with a five-frame delay, MOTA decreases by only 1.4%, whereas the random attack scenario results in a 60.5% drop.
{keyfinding}
Key Findings.  In contrast to 3D object detection tasks, MOT appears to rely more heavily on camera inputs. This may be attributed to the fact that MOT does not require precise 3D bounding boxes; instead, the rich texture information in camera images may offer more effective contrastive representations than sparse point clouds. As a result, DejaVu attacks can substantially impair MOT performance, particularly under Uni-DejaVu (camera delay) and Mul-DejaVu with random delay scenarios.
V-C Hardware-in-the-loop Testbed
In this part, we present our hardware-in-the-loop experiment for DejaVu, conducted on an AE testbed. As illustrated in Fig. 9, the testbed consists of three Raspberry Pi, representing camera, LiDAR, and fusion node. To ensure reproducibility and precise control over experimental conditions, we utilize KITTI Tracking Dataset for the respective camera and LiDAR node. To enable TSN and AE functionality, each Raspberry Pi is equipped with a RealTime TSN HAT and a media converter, and the nodes are connected via an AE switch. ROS2 serves as the middleware for data distribution (i.e., publishing/subscribing messages). Published messages contain timestamps and sensor content (.png file for camera and .bin file for LiDAR), which are sent sequentially at 1-frame per second. The fusion node subscribes to the camera and LiDAR topics and aligns the sensor messages according to their timestamps using ROS2’s $ApproximateTimeSynchronizer$ filtering utility. This utility maintains queues of incoming camera and LiDAR messages and aligns them based on their timestamps, permitting a slight delay between matched messages. The fusion node then performs MMF-based downstream processing on the aligned sensor data.
DejaVu Attack Impact on AE Testbed. In this experiment, we assume that the attacker is targeting the LiDAR sensor with the attacker capability $\scriptstyle C_{2}$ (see Section III-B1). As shown in Fig. 10(top), the attacker first publishes six LiDAR messages with benign timestamp; hence they are very close to the camera messages’ timestamps. The attacker then deliberately introduces a constant offset of 5 seconds to the timestamps of the next eleven LiDAR messages, while continuing to transmit genuine, real-time data. As a result, the timestamps of these LiDAR messages are abruptly increased, as depicted in the figure.
In the Fig. 10(bottom), the message content is accurately synchronized for camera and LiDAR pairs, for all benign timestamps, which are shown in green rectangles. When the attacker sends manipulated timestamps, the synchronizer forces camera messages to align with LiDAR messages inaccurately, which are shown using red arrows in the figure. This demonstrates that though the contents of camera and LiDAR messages were sequential, the synchronizer fails to align them accurately during the attack, as it prioritizes the timestamp of the messages. The impact of this attack is similar to the attack we see in Fig. 1(c).
Figure 13: Impact of delayed but valid LiDAR data on Autoware: the induced SLAM drift propagates to the planning module, ultimately causing the vehicle to collide with the curb.
V-D Simulation-Based End-to-End AD Setup
We conduct our experiments using Autoware, an open-source full-stack autonomous driving framework. Autoware is widely adopted in both commercial and public-sector deployments, including Level 4 autonomy trials and government-funded programs (e.g., the U.S. Department of Transportation’s CARMASM initiative). Its extensive use in real-world systems makes it a realistic and representative platform for evaluating the safety and robustness of autonomous driving pipelines. To simulate real-world driving scenarios in a reproducible and controllable setting, we integrate Autoware with AWSIM Lab—a Unity-based, open-source simulator developed as part of the Autoware ecosystem—that provides high-fidelity urban environments and realistic sensor simulation. The simulated vehicle is equipped with a representative sensor suite including GNSS, IMU, three Velodyne VLP-16 LiDARs, and a traffic light camera. ROS2 (Humble) works as the middleware, enabling seamless and real-time communication between AWSIM’s simulated sensors and Autoware full autonomy stack, allowing us to evaluate the impact of DejaVu attack in a safe yet realistic environment. Our experiments are conducted on Tokyo’s Nishishinjuku district road map.
DejaVu Attack Impact on AD Simulation. We demonstrate the DejaVu attack in a full-stack autonomous driving pipeline, targeting the LiDAR sensor under attacker capability $\scriptstyle C_{3}$ (see Section III-B1). Given LiDAR’s dominant role in 3D perception, the attacker impersonates three legitimate LiDAR nodes by subscribing to their respective ROS2 topics and monitoring inter-frame intervals (10 Hz) to predict the next transmission times. The attacker stores recent point cloud messages and, at the time of attack, publishes forged messages with previously captured data with updated timestamps—just before the expected legitimate message. This increases the likelihood that the forged message is selected by the time-based synchronizer if it aligns more closely with the timestamps of other modalities (e.g., IMU, camera). Fig. 11 illustrates how these delayed messages are positioned and transmitted on the same topics, effectively impersonating genuine LiDAR messages in real time.
When the forged messages are utilized in the downstream tasks, in the most severe case, the system completely misses the presence of an actual oncoming vehicle, resulting in a head-on collision at an intersection—despite nearby objects being within sensor range (Fig. 12a). This constitutes a false negative perception failure with life-threatening implications. In another scenario, delayed LiDAR data causes the ego vehicle to perceive a non-existent obstacle—a vehicle that has already passed—leading to unnecessary emergency braking (Fig. 12b). This false positive event can create rear-end collision risks. In both cases, the outdated sensor data was used to repeatedly overwrite fresh messages, degrading the temporal integrity of the perception pipeline.
Beyond object-level failures, we observe broader impacts across the autonomy stack. The tracker may assign separate IDs to genuine and delayed instances of the same object, interpreting them as distinct entities. Similarly, temporal inconsistencies introduced in the LiDAR stream desynchronize SLAM modules, leading to localization drift and control failures such as veering off-lane or collisions with curbs and roadside objects (as shown in Fig. 13).
VI Discussion
VI-A Attack Limitations
While the DejaVu attack demonstrates the vulnerability of multimodal perception systems to temporal misalignment, there are several limitations that constrain its applicability in real-world settings. First, the attack has not been validated on a physical vehicle platform, where additional practical challenges such as sensor noise, actuator delays, and system integration issues may affect both the feasibility and effectiveness of the attack. Besides, the attack assumes a high level of knowledge about the target system, including the sensors and network architecture. In practice, the attack requires adversarial access to the in-vehicle network. Although not trivial, prior work shows that physical access through the OBD-II port, remote exploitation of infotainment systems, or supply-chain insertion can provide such capabilities [18]. However, the adversary would still need to perform an exploration phase to gather this information, which introduces additional complexity and may limit the ability to execute the attack stealthily.
Moreover, the evaluation does not account for real-world network-induced delays and variability, which could impact the targeted timing manipulation strategies. In actual deployment, the presence of stochastic latency and jitter may reduce the precision with which an attacker can control temporal misalignment, potentially diminishing the attack’s effectiveness. Finally, the evaluation was conducted exclusively on an offline dataset with relatively low frame rates. Real-time systems typically operate at higher frame rates, and the metrics and thresholds used in this offline evaluation may not directly translate to real-world performance. Consequently, the practical impact of the attack on deployed autonomous systems may differ from the results observed in the experimental study.
VI-B Defense Strategies Against DejaVu
Hardening defenses can reduce the attack surface by securing sensor timestamps before fusion. This can be achieved via authenticated, hardware-anchored timestamps with cryptographic signatures, monotonic sequence numbers, and hardware-level clocks, such as network interface controller/SoC real-time clock (RTC). Combining multiple time sources—PTP, GNSS, and local RTCs—further strengthens temporal integrity. While these measures add overhead, they significantly raise the barrier to adversarial manipulation.
Detection defenses can identify temporal misalignments in real time. Techniques include intermodality temporal-consistency analysis of embeddings, kinematic cross-checks using IMU, odometry, and other controller area networks (CAN) signals, and statistical monitoring of monotonicity, jitter, and freshness counters, before fusion. These mechanisms can detect out-of-order or replayed frames, allowing rapid response to potential attacks.
Mitigation techniques can limit the impact of detected misalignments on vehicle control. Delay-aware adaptive fusion compensates for inter-modal lags, weighs sensor inputs, and computes confidence scores. If confidence is low, conservative measures, such as slowing down, increasing spacing, or lowering autonomy, can be applied, ensuring safety even under temporal attacks.
VII Conclusion
This work presents DejaVu, a temporal misalignment attack that exploits synchronization vulnerabilities in multimodal perception systems for autonomous driving. Through extensive evaluations on state-of-the-art 3D object detection and multi-object tracking models, we uncover modality-specific vulnerabilities: 3D detection models are predominantly reliant on LiDAR and suffer severe degradation—with up to 88.5% drop in mAP—from even a single-frame LiDAR delay, while MOT models exhibit heightened sensitivity to camera stream disruptions, with MOTA dropping by 73% under just three-frame camera delays. These findings highlight the critical need for synchronization-aware design in perception architectures and emphasize the importance of robust temporal consistency checks in safety-critical autonomous systems.
LLM Usage Disclosure.
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.